# Chatbot Arena Meets Nuggets: Towards Explanations and Diagnostics in the Evaluation of LLM Responses

### v1

@@ 4 Experimental Results
(e) Knowledge-intensive
(f) Reasoning-intensive
(g) Subjective
@@
Figure 5: Confusion matrices comparing human and nugget preferences across eight query classes from the Search Arena dataset. Threshold of 0.1 is used to treat the nugget preference as a tie.
4.1 Query Classification Analysis
In this analysis we use query classification to better understand the cases where nugget preferences and human preferences are not aligned. When the nugget scores and the human prefer opposite sides of a battle, we refer to this situation as a “preference inversion”, or simply inversion.
We suspect that inversions might vary across different types of queries. To investigate, we followed Rosset et al. (2024) but used the newer GPT4.1 to rate each query on a scale of 0–10 across eight different categories. Then, we classify each query into its maximum scoring category or categories (allowing for ties). To further strengthen the category signals, we exclude queries with a maximum score less than seven from this classification. Raw distributions of the query ratings per category and sample queries from each class are available in Section A.2.
@@ 4.1 Query Classification Analysis
(1) Multi-faceted 19% 298
(2) Ambiguous 19% 194
(3) Assumptive 18% 28
(4) Incompleteness 18% 627
(5) Knowledge-intensive 16% 1135
(6) Reasoning-intensive 16% 283
(7) Subjective 13% 598
(8) Harmful 5% 82
Table 1: Inversion percentages and query frequencies of eight different query classes in the Search Arena dataset.
As shown in Table 1, the portion of inversions for multi-faceted, ambiguous, assumptive, and incomplete queries is higher than that of subjective, knowledge-, and reasoning- intensive queries. This suggests that inversions are more likely when queries allow for multiple valid interpretations or are under-specified.
@@
Overall, the knowledge-intensive class shows the highest preference alignment—58.8% and 53.9% for modelA and modelB wins, respectively (see Figure 5). This finding suggests that nuggetization is most effective for research-y queries requiring retrieval augmentation.
Language Inversion (%) Query Count
(1) German 19% 240
(2) English 16% 3089
(3) Portuguese 16% 148
(4) Chinese 16% 324
(5) Russian 14% 459
(6) French 13% 150
(7) Others 15% 642
Table 2: Inversion percentages and query frequencies of the six most common languages in the Search Arena dataset.
@@
Figure 6: Confusion matrices comparing human and nugget preferences across six different languages that each account for at least 3% of the Search Arena dataset. Threshold of 0.1 is used to treat the nugget preference as a tie.
4.2 Query Language Analysis
We next analyzed AutoNuggetizer effectiveness across the six most frequent query languages, each representing at least 3% of the dataset. Previously, the tool had only been run on English responses, and there are likely to be language effects in the breakdown of inversions.
As shown in Table 2, German exhibits the highest inversion rate (19%), while French shows the lowest (13%). The confusion matrix for German (see Figure 6) reveals that it has the smallest portion of ties in human preferences, leading to more anti-diagonal disagreements. Limited human-voted ties suggest that the LLMs participating in the battles often differ in their ability to handle German queries. Additionally, assuming a similar distribution of query categories across languages, the higher inversion rate among German queries points to the AutoNuggetizer being less effective in this language as well. Due to the limited dataset size, we leave language-specific query classification analysis for future work.
@@ 5 Discussion
At a high level, we find that these results are quite strong, given that we have only examined one factor that might influence LLM response quality. For example, human preferences might be affected by how citations are presented, the fluency or organization of the responses, the presence of aids such as tables and headings, as well as a myriad of other factors. Nevertheless, with our automatically computed fact recall metric, we are able to predict human preferences over 50% of the time. This is quite remarkable in our opinion, and potentially points to the explanatory power of nugget scores.
Although we have only begun explorations, from here it is possible to imagine how our approach can be extended into providing diagnostic information to system builders. Missing nuggets can be attributed to different causes: a relevant document that was not retrieved or the LLM ignoring relevant context present in the prompt. Different failure modes would lead to different courses of action. For example, a retrieval failure could point to the need for improvements to the embedding model, and perhaps the battle result can be adapted into an additional training example.
Building on these initial insights, this paper serves as the first stage of exploring nugget evaluation for search-based arena battles. In our preliminary analyses, we missed out on evaluating a few things that we keep as future work:
First, we would like to consider an LLM-as-a-judge evaluation and check correlations between an LLM assessor and our nugget evaluation. Next, we currently filter to include only single-turn conversations; in the future we would like to evaluate multi-turn conversations whenever per-turn user votes become available. Finally, battles in the Search Arena dataset include URLs to source documents as well as web search traces such as search queries. This information can be valuable for generating nuggets and assessing the factuality of LLM-generated responses. We leave as future work the exploration of leveraging grounded URLs to automatically generate nuggets based on the retrieved content.
@@ 6 Conclusion
This work explores the use of nugget-based evaluation to assess large language model (LLM) competitions in Search Arena, a benchmark for side-by-side comparisons of search-augmented model responses. By generating and scoring atomic facts (nuggets), we offer a more interpretable and diagnostic alternative to traditional human preference-based evaluations.
Our results show strong alignment between nugget-based preferences and human judgments, particularly for knowledge-intensive queries. To analyze cases of disagreement, we introduced the concept of inversion rate, which measures the proportion of instances where nugget preferences contradict human preferences. Higher inversion rates were found in multi-faceted, ambiguous, and assumptive queries, suggesting these query types are more challenging for automated evaluation. Additionally, language-level analysis reveals that German queries have the highest inversion rate among the major languages, pointing to potential limitations in nuggetization quality for certain languages.
@@ A.1 Dataset Statistics
Figure 8: Histogram showing the classified attributes for 5,103 single-turn queries in Search Arena. We use GPT4.1 with prompt from Researchy Questions Rosset et al. (2024) to output a score between 1–10 for each attribute.

### v2

@@ 4.1 Query Classification Analysis
(e) Subjective
(f) Knowledge-intensive
(g) Reasoning-intensive
@@
Figure 5: Confusion matrices comparing human and nugget preferences across eight query classes from the Search Arena dataset. A threshold of 0.07 is used to treat nugget preference scores as a tie.
Category Inversion (%) Query Count
(1) Ambiguous 19% 196
(2) Assumptive 18% 28
(3) Multi-faceted 18% 299
(4) Incompleteness 16% 631
(5) Subjective 15% 601
(6) Knowledge-int. 15% 1142
(7) Reasoning-int. 14% 288
(8) Harmful 9% 92
@@
(1) German 20% 244
(2) English 17% 3117
(3) Chinese 16% 328
(4) Portuguese 16% 150
(5) Russian 15% 460
(6) French 13% 151
(7) Others 16% 647
Table 1: Inversion percentages and query frequencies across (a) different query categories and (b) languages in the Search Arena dataset.
As shown in Table 1 (side a), the portion of inversions for ambiguous, assumptive, multi-faceted, and incomplete queries is higher than that of subjective, knowledge-, and reasoning-intensive queries. This suggests that inversions are more likely when queries allow for multiple valid interpretations or are underspecified.
@@
Overall, the knowledge-intensive class shows the highest preference alignment—58.8% and 55.7% for modelA and modelB wins, respectively (see Figure 5). This finding suggests that nuggetization is most effective for researchy queries requiring retrieval augmentation.
4.2 Query Language Analysis
@@
Figure 6: Confusion matrices comparing human and nugget preferences across six different languages that each account for at least 3% of the Search Arena dataset. A threshold of 0.07 is applied to treat nugget preference scores as a tie.
Next, we analyzed the AutoNuggetizer effectiveness across the six most frequent query languages, each representing at least 3% of the dataset. Previously, the AutoNuggetizer had only been run on English responses, and there are likely to be language effects in the breakdown of inversions.
As shown in Table 1 (side b), German exhibits the highest inversion rate (20%), while French shows the lowest (13%). The confusion matrix for German (see Figure 6) reveals that it has the smallest portion of ties in human preferences, leading to more anti-diagonal disagreements. Limited human-voted ties suggest that the LLMs participating in the battles often differ in their ability to handle German queries. Additionally, assuming a similar distribution of query categories across languages, the higher inversion rate among German queries points to the AutoNuggetizer being less effective in this language as well. Due to the limited dataset size, we leave language-specific query classification analysis for future work.
4.3 Nugget Generation without URL Contents
Figure 7: Confusion matrix comparing human and nugget preferences, using LLM responses as the sole source for nugget generation. A threshold of 0.1 is applied to treat nugget preference scores as a tie.
To assess the impact of scraped URL contents on the effectiveness of nugget-based evaluations, we generate nuggets using only the two LLM responses in this study. Out of the 5,103 single-turn battles, 51 were excluded due to nugget generation failures. As shown in Figure 7, the effectiveness of nugget generation using only LLM responses is comparable to that of using both URL contents and LLM responses. Specifically, the agreement with human preferences when modelA is the winner is 54.8%, versus 54.7% when URL contents are included. For modelB as the winner, the agreement is 52.4%, compared to 52.5% with URL contents. Due to the smaller number of nuggets generated from LLM responses alone, the resulting nugget scores are more discrete. Consequently, we use a threshold of 0.1 to classify ties, instead of 0.07 as used in the case of nugget generation with URL contents.
These results suggest that LLM responses alone can serve as a viable source for nugget generation when external evidence is unavailable or unreliable. However, incorporating URL contents may still be beneficial in increasing nugget diversity and grounding, especially in cases requiring factual precision.
4.4 LLM-as-a-Judge Evaluation
To analyze the correlation between human and LLM preferences, we experiment with GPT4.1 as a judge. We modify the chain-of-thought prompt provided in Rackauckas et al. [25] (refer Appendix B). For each evaluation, we provide the user query along with the two model responses—randomly ordered to mitigate positional bias—as input to GPT4.1. The model is instructed to output its reasoning and final verdict in a structured JSON format.
Figure 8: Confusion matrix comparing human and GPT4.1 preferences.
Figure 8 presents the confusion matrix comparing human preferences with those of the GPT4.1 judge. Compared to the nugget-based evaluation in Figure 4, we observe a stronger alignment between GPT4.1 and human judgments for clear winners: 1,137 vs. 938 agreements for modelA, and 1,161 vs. 924 for modelB. However, GPT4.1 struggles significantly with identifying ties—including cases where both responses are poor—labeling only 4.25% of the single-turn queries as such. This narrow margin for tie predictions leads to a higher frequency of preference inversions when using LLM-as-a-judge, with 1,102 inversions compared to 817 under the nugget-based evaluation. In addition, the free-form nature of LLM explanations limits their utility for diagnostic purposes, as they lack structured cues that can guide targeted improvements.
@@ 5 Discussion
At a high level, we find that these results are quite strong, given that we have only examined one factor that might influence LLM response quality. For example, human preferences might be affected by how citations are presented, the fluency or organization of the responses, the presence of aids such as tables and headings, as well as a myriad of other factors. Nevertheless, with our automatically computed fact recall metric, we can predict human preferences over 50% of the time. This is quite remarkable in our opinion, and potentially points to the explanatory power of nugget scores.
Although we have only begun explorations, from here it is possible to imagine how our approach can be extended into providing diagnostic information to system builders. Missing nuggets can be attributed to different causes: a relevant document that was not retrieved or the LLM ignoring relevant context in the prompt. Different failure modes would lead to different courses of action. For example, a retrieval failure could point to the need for improvements to the embedding model, and perhaps the battle result can be adapted into an additional training example. Building on these initial insights, this paper serves as the first stage of exploring nugget evaluation for search-based arena battles.
6 Limitations
Our current evaluation focuses exclusively on single-turn conversations, as the Search Arena dataset lacks per-turn user judgments for multi-turn interactions. Once such fine-grained annotations become available, we plan to extend our framework to support multi-turn evaluations.
While battles in the dataset include URLs to web search results—which are valuable for grounding and factuality assessment—there are key limitations. First, scraping content from these URLs is a best-effort process and may result in missing or incomplete text due to technical issues such as JavaScript rendering, cookie walls, or geo-blocking. Second, web content is dynamic; the scraped content may not reflect what the LLM originally accessed when generating its response. To improve reproducibility, we recommend that future dataset releases include archived snapshots of the referenced URLs.
Lastly, in this study, we used a specific model for nugget generation. Exploring the impact of different models on nugget quality and agreement remains an open direction for future work, particularly to assess consistency across different LLMs as nuggetizers.
@@ 7 Conclusion
This work explores nugget-based evaluation to assess large language model (LLM) competitions in Search Arena, a benchmark for side-by-side comparisons of search-augmented model responses. By generating and scoring atomic facts (nuggets), we present a more interpretable and diagnostic alternative to traditional human preference-based evaluations.
Our results demonstrate a strong alignment between nugget-based preferences and human judgments, especially for knowledge-intensive queries. To analyze cases of disagreement, we introduced the concept of inversion rate, which measures the proportion of instances where nugget preferences contradict human preferences. Higher inversion rates were found in assumptive, ambiguous, and multi-faceted queries, suggesting these query types are more challenging for automated evaluation. Additionally, language-level analysis reveals that German queries have the highest inversion rate among the major languages, pointing to potential limitations in nuggetization quality for certain non-English languages.
We further showed that nuggetization using only LLM responses—without access to retrieved URL contents—remains highly effective, with human agreement levels nearly identical to those obtained when URL content is included. This robustness demonstrates the practicality of nugget-based evaluation, even in retrieval-limited settings. We also evaluated an LLM-as-a-judge baseline using GPT4.1 with chain-of-thought prompting. While it exhibited higher agreement with human preferences in clear win/loss cases, it struggled with identifying ties, labeling only 4.25% of queries as such. Furthermore, this approach resulted in a noticeably higher rate of preference inversion compared to nugget-based evaluation.
@@
8 Broader Impact
By leveraging a nugget-based evaluation to explain human preferences and diagnose model behavior, our work contributes to the development of more transparent and accountable LLMs. This contributes to more trustworthy AI systems, especially in high-stakes domains where factual accuracy and interpretability are critical. By enabling diagnostic insights, our work supports the development of more reliable and equitable AI tools, helping mitigate misinformation and promoting informed decision-making across society.
@@ A.1 Dataset Statistics
(a) Assumptive
(b) Multi-faceted
(c) Incompleteness
(d) Ambiguous
(e) Subjective
(f) Knowledge-Intensive
(g) Reasoning-Intensive
(h) Harmful
Figure 10: Histogram showing the classified attributes for 5,103 single-turn queries in the Search Arena dataset. We use GPT4.1 with prompt from Researchy Questions [26] to output a score between 0–10 for each attribute.
@@ A.2 Query Classification
Appendix B GPT4.1 Judge Prompt Details
{mdframed}
[font=, roundcorner=10pt, linewidth=1pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=5pt, innerbottommargin=5pt] Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants tasked to answer the question displayed below. You should choose the assistant that best answers the user question.
Your evaluation should consider factors such as the correctness, helpfulness, completeness, accuracy, depth, and level of detail of their responses. Details are only useful if they answer the user question. If an answer contains non-relevant details, it should not be preferred over one that only use relevant information.
Begin your evaluation by explaining why each answer correctly answers the user question. Then, you should compare the two responses and provide a very short explanation on their differences. Avoid any position biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Be as objective as possible. Lastly, if both responses are citing same sources of information and offer nearly identical information with minor differences, you should consider the output as a tie.
After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is better, "[[B]]" if assistant B is better, and "[[Tie]]" for a tie.
[The Start of User’s Question]
{$query$}
[The End of User’s Question]
[The Start of Assistant A’s Answer]
{$answer_{a}$}
[The End of Assistant A’s Answer]
[The Start of Assistant B’s Answer]
{$answer_{b}$}
[The End of Assistant B’s Answer]
Figure 11: Prompt used by GPT4.1 judge to evaluate the model answers in Search Arena.
Figure 11 illustrates the chain-of-thought prompt modified and referenced originally from RAGElo [25]. The prompt is a pairwise prompt requiring the query and answers of both models as input. Next, GPT4.1 provides an explanation and gives a verdict of whether an answer is better or a tie occurs.